Circular No. 7408

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000bk in NGC 4520 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a spectrum of SN 2000bk (cf. IAUC 7402), taken on Apr. 25.19 UT by M. Calkins with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova after maximum light. The spectrum features well-developed lines of Si II, Na I, and Fe II, and narrow interstellar Na I D absorption (equivalent width 0.09 nm) from the host galaxy is also present. The NASA/IPAC Extragalactic Database (NED) recession velocity for NGC 4520 is 7628 km/s, implying a supernova expansion velocity of 9800 km/s for the Si II (rest 635.5 nm) line. The age of the supernova based on its spectral features is 8 ± 3 days past maximum light (Riess et al. 1997, A.J. 114, 722).

Preliminary photometry of SN 2000bk by K. Krisciunas, University of Washington; and R. McMillan, Apache Point Observatory (APO), using the APO 3.5-m telescope and 5′′.6-diameter software aperture (estimated uncertainties are ± 0.05 for the V magnitude and ± 0.03 for the colors): Apr. 24.19 UT, V = 17.24, B −V = +0.78, V −R = +0.01, V −I = −0.20.

SUPERNOVA 2000bs IN UGC 10710 Jha et al. also report that a spectrum of SN 2000bs (cf. IAUC 7406),

obtained by Calkins as above on Apr. 25.45 UT, shows it to be a type-II supernova at an early epoch. The spectrum exhibits a very blue continuum with prominent, broad Hβ absorption and a hint of broad Hα emission. The supernova photospheric expansion velocity (measured using the Hβ profile) is 11 900 km/s, adopting the NED recession velocity of 8387 km/s for the host galaxy.

Preliminary photometry of SN 2000bs by Krisciunas and McMillan, obtained as above: Apr. 24.24 UT, V = 18.10, B − V = +0.19, V − R = +0.19, V − I = +0.40.

COMET P/2000 G1 (LINEAR) The following improved orbital elements from MPC 39792 confirm that

this comet (cf. IAUC 7396) is of short period:

T = 2000 Mar. 9.796 TT  ω = 343.^{o}278
e = 0.67397  Ω = 191.048 }2000.0
q = 1.00293 AU  i = 10.401
a = 3.07620 AU n^{o} = 0.182676 P = 5.40 years
2000 April 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green